How many grams of ammonium chloride must be added to 0.050 kg of water to produce a
solution that freezes at -5.00 °C? The cryoscopic constant of water is 1.86 °C kg mol-1
. The
molar mass of ammonium chloride is 53.5 g mol-1
1
Expert's answer
2020-06-09T07:20:04-0400

ΔT = K ×m

ΔT is change in freezing point

K is cryoscopic constant

m is molality of the solution

Here

ΔT = 5

K= 1.86

m = ΔT / K

= 5/ 1.86

=2.69

suppose x gm of Ammonium Chloride is required

m = ( x/ 53.5)/ 0.05

2.69 = (x/ 53.5)/ 0.05

x = 2.69 × 0.05 × 53.5

= 7.2 g